# High Dimensional Ensemble Kalman Filter

### v1

@@ Abstract
The ensemble Kalman Filter (EnKF), as a fundamental data assimilation approach, has been widely used in many fields of earth science, engineering and beyond. However, there are several unknown theoretical aspects of the EnKF, especially when the state variable is of high dimensional accompanied with high resolution observation and physical models. This paper first proposes several high dimensional EnKF methods which provide consistent estimators for the important forecast error covariance and the Kalman gain matrix. It then studies the theoretical properties of the EnKF under both the fixed and high dimensional state variables, which provides the mean square errors of the analysis states to the underlying oracle states offered by the Kalman filter and gives the much needed insight into the roles played by forecast error covariance on the accuracy of the EnKF. The accuracy of the data assimilation under the misspecified physical model is also considered. Numerical studies on the Lorenz-96 and the Shallow Water Equation models illustrate that the proposed high dimensional EnKF algorithms perform better than the standard EnKF methods as they provide more robust and accurate assimilated results. The high dimensional EnKF is applied to assimilate sea temperature in Northwest Pacific, which showed more accurate out-sample performance than the existing methods.
Data Assimilation,
Ensemble Kalman Filter,
High Dimension Imputation,
keywords:
[…]
, and
@@ 1 Introduction
Kalman Filter (Kalman, 1960) is a fundamental approach for forecasting and data assimilation for a partially observed dynamic system. Kalman Filter and its later developments (Ensembled Kalman Filter, three and four–dimensional variational assimilation (3D-Var, 4D-Var) (Pailleux, 1990; Courtier et al., 1998) are based on the optimal estimation theory under the multivariate normal distribution. Data assimilation is a statistical procedure that revises a forecast based on the physical model to the state variables of the dynamic system upon receiving new observations. It combines information from numerical physical model forecasts with observed data to obtain the best possible description of a dynamical system and its uncertainty (Evensen, Vossepoel and van Leeuwen, 2022). The assimilated state variables at a time are used to generate the forecast for the next time step. For chaotic systems sensitive to initial values, data assimilation can significantly improve the accuracy of the forecast by using optimized state variables as initial values (Lorenz, 1963).
Although much applied in numerical prediction for weather and oceanography, Kalman Filter and its later versions are the methodological core for scientific algorithms over a range of fields, including hydrology, remote sensing data utilization, air quality prediction, carbon data assimilation, robotic control and automatic driving. Data assimilation methods also provide the underlying algorithms and predictive techniques for building digital twins for social and economic systems.
@@
The limitations of the Kalman Filter are (i) it assumes a linear model system which is known to be restrictive for many non-linear dynamic systems; and (ii) the forecast error covariance matrix and Kalman Gain matrix are expensive to store and update when the dimension of the state vector is large. The Ensemble Kalman Filter (EnKF; Evensen, 1994) was proposed to overcome these limitations, where the forecast error covariance matrix is estimated by the sample covariance based on generated forecast ensembles. The EnKF only requires storing and operating on reduced size state vectors and the ensembles can be generated parallelly. These advantages have made the EnKF an effective data assimilation scheme and has been widely used in atmospheric and oceanic sciences, and engineering.
Estimating the forecast error covariance matrices is a key in the Kalman Filter and the EnKF schemes, which is made in the EnKF by the sample covariance matrix of the forecast ensembles. Past works on the EnKF found that the sampling error can lead to underestimation of the forecast error covariance and can eventually cause the filter to diverge (Anderson and Anderson, 1999; Constantinescu et al., 2007). The filter divergence refers to a decreasing ability to correct the ensemble state towards the observations after certain rounds of assimilation. One reason for the filter divergence is the variance of the ensemble forecast error becomes too small when there are unaccounted model errors (Lorenc, 2003); the other is due to the substantial mismatch between the estimated and true error covariance matrices (Houtekamer and Mitchell, 1998; Constantinescu et al., 2007).
@@
Along with advances in remote sensing technology in measuring the earth system, high spatial resolution observations of the earth system are increasingly available, which leads to high dimensional geophysical models with high dimensional state vectors (Stroud et al., 2010; Guinehut et al., 2012; Lellouche et al., 2013). High resolution data assimilation and forecasting techniques are much needed in modern meteorology and atmospheric environment systems. For example, in global CO2 calculation, the dimension of the state variable can reach tens of millions. The high dimensional state vectors lead to high dimensional  forecast error covariance and Kalman Gain matrices. The high dimensionality is known to cause the non-convergence of the forecast error covariance matrix (Bai and Yin, 1993). In practice, the high dimensional effect is felt in the form of the filter divergence and unreliable data assimilation. Geophysicists have proposed two methods to counter the problems. One is the localization method (Furrer, Genton and Nychka, 2006; Furrer and Bengtsson, 2007) that sets covariances to zero if the geographical distances between two sites are bigger than a threshold; and the other is the inflation method that enlarges the forecast error covariance by adding either a multiplicative factor or an additive term to avoid filter divergence. However, these methods lack support of statistical theory.
The last two decades have witnessed rapid development in high dimensional statistical inference. Consistent high dimensional  covariance estimators have been proposed, which include the banding, tapering and thresholding estimators (Bickel and Levina, 2008a, b; Cai, Zhang and Zhou, 2010). Both the banding and tapering estimators require a bandwidth parameter, and the thresholding estimator needs the threshold level. Qiu and Chen (2015) proposed bandwidth and thresholding level selection methods by minimizing an empirical estimate of the expected squared Frobenius norms of the estimation error matrix.
This paper proposes several high dimensional EnKF methods that utilize the high dimensional covariance estimators, and conducts a theoretical analysis of the proposed high dimensional EnKF. We also provide a theoretical analysis of the performance of the EnKF under the fixed dimensional setting since such a study appears lacking and is important. It is shown that with the use of the high dimensional covariance matrix estimators, consistent estimation of the Kalman Gain matrix and the underlying analysis states can be attained by the proposed high dimensional EnKF procedure. Specifically, we establish the mean square error (MSE) bounds of the analysis states from the EnKF methods to the underlying analysis under both fixed and high dimensional cases, as well as for both correct and imperfect physical models.
@@
Our analysis shows that consistent estimator of the forecast error covariance leads to the consistency of the Kalman Gain matrix by establishing an inequality that bounds the error of estimation of the Kalman Gain matrix by that of the forecast error covariance, which implies that, even at the high dimensional setting, the MSEs of the analysis states by the EnKF could be controlled by the error bounds of the high dimensional forecast error covariance estimators. Thus, consistent high dimensional covariance estimators would make the high dimensional EnKF consistent under certain conditions. Indeed, our theoretical studies show that both the forecast and analysis states of the EnKF converge to their underlying counterparts of the Kalman Filter if the forecast error covariance can be estimated consistently.
The investigation of the high dimensional EnKF under the misspecified model gives a more realistic description of the performance of the EnKF. Under the misspecified model, the MSE not only depends on the estimation error of the forecast error covariance matrices but also the discrepancy between the misspecified and the true model. Simulation studies on the Lorenz-96 model and the Shallow Water Equation model suggest that the proposed high dimensional EnKF algorithm could achieve more accurate and robust assimilation results than the widely used inflation-based EnKF method. The application to the sea temperature data assimilation in Northwest Pacific provides a comprehensive and accurate description of the spatiotemporal distribution of the three-dimensional sea temperature field by combining the ocean model with satellite data, which also illustrates the superiority and usefulness of the proposed high dimensional EnKF. This means the proposed high dimensional EnKF has wide application prospects in earth science like global carbon calculation, weather forecasting and the production of high-resolution data sets.
@@
The paper is structured as follows. Section 2 reviews the Kalman Filter and the EnKF. Theoretical properties of the EnKF for the fixed dimensional case are given in Section 3. Section 4 presents several consistent estimators of the high dimensional forecast error covariance and the key Kalman gain matrix, and establishes the theoretical properties of the EnKF for high dimensional case. The effects of imperfect or misspecified models on the accuracy of the EnKF are provided in Section 5. Results of empirical data assimilation based on two well known geophysical models are reported in Section 6 and the application to the sea temperature data assimilation is presented in Section 7 , followed by a conclusion in Section 8. Details of the proposed high dimensional EnKF algorithms, proofs of the theoretical results and additional results are given in the Supplementary Material (SM).
@@ 2 Review on Kalman and Ensemble Kalman Filters
where $\mathbf{x}_{t}$ is the $p-$dimensional not fully observed state vector of a dynamic system at time $t$, $\mathbf{M}_{t}$ is the linear model operator (also called evolution or forecast operator), $\mathbf{w}_{t}$ is the error term that follows a zero-mean multivariate normal distribution with $\mathbf{Q}_{t}$ as the covariance, $\mathbf{y}_{t}$ is the $q-$dimensional observation vector at time $t$, $\mathbf{H}_{t}$ is a linear observation operator which relates the state $\mathbf{x}_{t}$ to the observation $\mathbf{y}_{t}$, and $\boldsymbol{\varepsilon}_{t}$ is the observation error with $\mathbf{R}_{t}$ being its covariance matrix.
@@
In the update (assimilation) step, the goal is to update $\mathbf{x}_{t}^{f}$ according to the new observation $\mathbf{y}_{t}$ and to reduce the difference between the true state $\mathbf{x}_{t}$ and the forecast state $\mathbf{x}_{t}^{f}$. The joint distribution of $\left(\mathbf{x}_{t},\mathbf{y}_{t}\right)$ given $\mathbf{y}_{1:t-1}$ is
$\left(\begin{array}[]{c}\mathbf{x}_{t}\\ \mathbf{y}_{t}\end{array}\right)\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p+q}\left(\left(\begin{array}[]{c}\mathbf{x}_{t}^{f}\\ \mathbf{H}_{t}\mathbf{x}_{t}^{f}\end{array}\right),\left(\begin{array}[]{cc}\boldsymbol{\Sigma}_{t}^{f}&\boldsymbol{\Sigma}_{t}^{f}\mathbf{H}_{t}^{T}\\ \mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}&\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}\mathbf{H}_{t}^{T}+\mathbf{R}_{t}\end{array}\right)\right),$ (2.4)
which implies that
$\mathbf{x}_{t}\mid\mathbf{y}_{1:t}\sim{\mathrm{N}}_{p}\left({\mathbf{x}}_{t}^{a},{\boldsymbol{\Sigma}}_{t}^{a}\right)$ (2.5)
@@
which are, respectively, the analysis (assimilation) state and the associated analysis covariance matrix, and the Kalman gain matrix. It is noted that the analysis state $\mathbf{x}_{t}^{a}$ is the minimum variance unbiased estimator for the mean of $\mathbf{x}_{t}$ under the linear state-space model (2.1).
However, models in earth science or engineering are often nonlinear, which implies that the evolution model in (2.1) may not be linear so that $\mathbf{x}_{t}=\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}$, where $\mathcal{M}_{t}$ is a nonlinear model operator that can invalidate the forecast distribution in (2.2). Besides, for large dynamic systems, the dimension $p$ of the state vector $\mathbf{x}_{t}$ can be large so that the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ is expensive to compute and store. To tackle these problems, Evensen (1994) proposed the EnKF, where the forecast error variance matrix is no longer updated as (2.3) but by the sample covariance based on $n$ forecast ensembles. It hence only needs to store and operate on $n$ vectors of dimension $p$, and is convenient for parallel computing, and thus improves the computing efficiency and reduces the storage cost.
@@
where $\mathcal{M}_{t}$ is the possibly nonlinear model operator, $\mathbf{H}_{t}$ is a linear observation operator and $\boldsymbol{\varepsilon}_{t}$ is the observation error with $\mathbf{R}_{t}$ being its covariance matrix. It further assumes that the forecast error $\boldsymbol{\eta}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$ follows a Gaussian distribution given $\mathbf{y}_{1:t-1}$, that is $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}_{p}(\mathbf{0},\boldsymbol{\Sigma}_{t}^{f})$, where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)$. Here $\mathcal{M}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are assumed to be known for the time being as in the Kalman Filter.
@@
    Estimate the forecast error covariance matrix $\Sigma_{t}^{f}$ by the sample covariance matrix of the forecast ensemble, namely
@@
    and obtain an estimate of the Kalman gain matrix $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$. Here we note computationally efficient expressions
    $\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}$ $=(n-1)^{-1}\sum_{j=1}^{n}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\{}\mathbf{H}_{t}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\}}^{\mathrm{T}}\,\text{and}$ (2.11)
    $\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}$ $=(n-1)^{-1}\sum_{j=1}^{n}\big{\{}\mathbf{H}_{t}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\}}\big{\{}\mathbf{H}_{t}\big{(}\mathbf{x}_{t,j}^{f}-\hat{\mathbf{x}}_{t}^{f}\big{)}\big{\}}^{\mathrm{T}}$
    since the dimension $q$ of the observation is much smaller than that ($p$) of the state vector.
@@
In practice, the algorithm is initialized at time $t=0$ by drawing $\{\mathbf{x}_{0,j}^{a}\}_{j=1}^{n}$ independently from ${\mathrm{N}}_{p}\left(\mathbf{x}_{0}^{a},\Sigma_{0}^{a}\right)$ where $\mathbf{x}_{0}^{a}$ and $\Sigma_{0}^{a}$ are either chosen based on the physical model or by pre-running the physical model for a period of time. Then, the ensemble is propagated forward through time, alternating between the forecast and the analysis steps, according to the three-step procedure above. If $\mathbf{R}_{t}$ and $\mathbf{Q}_{t}$ are unknown, their estimates can be entertained to substitute them in (2.9) and (2.12).
The EnKF does not require the expensive recursions (2.3) and (2.6)-(2.7) used in the Kalman Filter. Instead, it estimates the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance or the reduced dimension counterpart (2.11) via the observation operator $\mathbf{H}_{t}$.
For the case with non-linear model operator as in (2.8), we can still define the underlying analysis state $\mathbf{x}_{t}^{a}$ by (2.6)-(2.7), where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)=\mathbb{E}(\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}\mid\mathbf{y}_{1:t-1})$ is the forecast state and $\boldsymbol{\Sigma}_{t}^{f}$ is the forecast error covariance defined by $\boldsymbol{\Sigma}_{t}^{f}=\operatorname{var}((\mathbf{x}_{t}-\mathbf{x}_{t}^{f})\mid\mathbf{y}_{1:t-1})$. In fact, $\mathbf{x}_{t}^{a}$ defined by (2.6)-(2.7) is the solution to the minimizing problem (2.13) with $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ replaced by $\mathbf{x}_{t}^{f}$ and $\boldsymbol{\Sigma}_{t}^{f}$. For the ease of reference, and to avoid differentiating between the linear and the non-linear cases, we also call $\mathbf{x}_{t}^{a}$ as the analysis state of the Kalman filter even in the non-linear model case. The next two sections will establish the convergence of the EnKF analysis state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman filter.
It is noted that the update formula (2.12) can be obtained by minimizing
$J(\mathbf{x})=\frac{1}{2}\left\{(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})^{\mathrm{T}}(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})+\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)^{\mathrm{T}}\mathbf{R}_{t}^{-1}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)\right\}$ (2.13)
with respect to $\mathbf{x}$. To appreciate this, note that by setting $\partial J(\mathbf{x})/{\partial\mathbf{x}}=\mathbf{0}$ and by the Sherman-Morrison-Woodbury formula, we get
$\mathbf{x}$ =$=\left\{\hat{\boldsymbol{\Sigma}}_{t}^{f}-\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{R}_{t}+\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{)}^{-1}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\right\}\left\{(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}\mathbf{\hat{x}}_{t}^{f}+\mathbf{H}_{t}^{\mathrm{T}}\mathbf{R}_{t}^{-1}\mathbf{y}_{t}\right\}$
$=\mathbf{\hat{x}}_{t}^{f}+\mathbf{\hat{K}}_{t}\big{(}\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}_{t}^{f}\big{)},$
which is the data assimilation formula (2.12). The objective function (2.13) is related to that of the 3D-Var or 4D-Var, but not the same. However, over the same time interval, under the assumption of perfect model with the same inputs (initial mean and covariance matrix, observation error covariance), the 4D-Var analysis at the end is equal to the analysis of the Kalman filter; see Lorenc (2003), Hamill (2006) and Constantinescu et al. (2007) for details.
3 Consistency for fixed dimension case
We first provide the upper bound of the estimation error of the estimated Kalman gain matrix and then establish the convergence of the EnKF analysis states to the underlying Kalman Filter under both the linear and the non-linear model operators for the fixed dimensional state vectors with a convergence rate. Indeed, it appears that such results are lacking in the literature. Additionally, the error bounds could further be extended to the high dimension cases, which will be discussed later in Section 4.

### v2

# aoas/sm_enkf

@@ Abstract
The ensemble Kalman Filter (EnKF), as a fundamental data assimilation approach, has been widely used in many fields of earth science, engineering and beyond. However, several theoretical aspects of the EnKF remain unknown, especially when the state variable is of high dimensional accompanied with high resolution observation and physical models. This paper first proposes several high dimensional EnKF (HD-EnKF) methods that provide consistent estimators for the important forecast error covariance and the Kalman gain matrix. It then studies the theoretical properties of the EnKF under both the fixed and high dimensional state variables, which provides the mean square errors of the analysis states to the underlying oracle states offered by the Kalman filter and gives the much needed insight into the roles played by the forecast error covariance on the accuracy of the EnKF. The accuracy of the data assimilation under the misspecified physical model is also considered. Numerical studies on the Lorenz-96 and the Shallow Water Equation models illustrate that the proposed HD-EnKF algorithms perform better than the standard EnKF methods as they provide more robust and accurate assimilated results. The HD-EnKF was applied to assimilate sea temperature in the Northwest Pacific, which showed more accurate out-sample performances than the existing methods.
Keywords: Data Assimilation; Dynamic System; Ensemble Kalman Filter; High Dimensional Imputation.
@@ 1 Introduction
Kalman Filter (Kalman, 1960) is a fundamental approach for forecasting and data assimilation for a partially observed dynamic system. Data assimilation is a statistical procedure that combines information from numerical physical model forecasts with observed data to obtain the best possible description of a dynamical system and its uncertainty (Evensen et al., 2022). The assimilated state variables at a time are used to generate the forecast for the next time step. For chaotic systems sensitive to initial values, data assimilation can significantly improve the accuracy of the forecast by using optimized state variables as initial values (Lorenz, 1963). Although much applied in numerical weather and oceanography prediction, Kalman Filter and its later developments (Ensemble Kalman Filter, three and four-dimensional variational assimilation (3D-Var, 4D-Var; Pailleux, 1990; Courtier et al., 1998)) are the methodological core for scientific algorithms over a range of fields, including hydrology, remote sensing data utilization, carbon data assimilation, robotic control and automatic driving. Data assimilation methods also provide the underlying algorithms and predictive techniques for building digital twins for social and economic systems.
@@
The limitations of the Kalman Filter are (i) it assumes a linear model which is restrictive for many non-linear dynamic systems; and (ii) the forecast error covariance and Kalman gain matrices are expensive to store and update when the dimension of the state vector is large. To overcome these limitations, Evensen (1994) proposed the Ensemble Kalman Filter(EnKF), where the forecast error covariance is estimated by the sample covariance based on generated forecast ensembles. The EnKF only requires storing and operating on much reduced-size state vectors and the ensembles can be generated in parallel. These advantages have made the EnKF an effective and widely used data assimilation scheme.
Estimating the forecast error covariance is a key in the EnKF, which is made in the EnKF by the sample covariance of the forecast ensembles. Past works on the EnKF found that the sampling error can lead to underestimation of the forecast error covariance and can eventually cause filter divergence(Anderson and Anderson, 1999; Constantinescu et al., 2007). Filter divergence refers to a decreasing ability to correct the ensemble state towards the observations after certain rounds of assimilation. High resolution data assimilation and forecasting techniques are much needed in modern meteorology and atmospheric systems like global carbon calculation and ocean data assimilation, which leads to high dimensional geophysical models with high dimensional state vectors. Besides, along with advances in remote sensing technology in measuring the earth system, there is a growing availability of high spatial resolution satellite observations (Stroud et al., 2010; Guinehut et al., 2012).
@@
The high dimensional state vectors lead to high dimensional  forecast error covariance and Kalman gain matrices. The high dimensionality is known to cause the non-convergence of the sample forecast error covariance (Bai and Yin, 1993), which is felt in the form of filter divergence and unreliable data assimilation in practice. Geophysicists have proposed two methods to counter the problems. One is the localization method (Furrer et al., 2006; Furrer and Bengtsson, 2007) that sets covariances to zero if the geographical distances are bigger than a threshold; and the other is the inflation method that enlarges the forecast error covariance by adding either a multiplicative factor or an additive term (Wang and Bishop, 2003; Liang et al., 2012; Constantinescu et al., 2007). However, the statistical properties of these EnKF methods are generally lacking, despite their widespread applications. Katzfuss et al. (2020) proposed Gibbs Ensemble Kalman Filter and Smoother, and particle Ensemble Kalman Filter and Smoother algorithms for high dimensional hierarchical dynamic models from a Bayesian perspective.
@@
This paper proposes several high dimensional EnKF (HD-EnKF) methods based on consistent high dimensional covariance estimators and conducts a theoretical analysis of assimilated state vectors. It is shown that consistent estimation of the forecast error covariance leads to the consistency of the Kalman gain matrix by establishing an inequality that bounds the error of estimation of the Kalman Gain matrix by that of the forecast error covariance. This implies that, even at the high dimensional setting, the mean square errors (MSEs) of the analysis states by the EnKF could be controlled by the error bounds of the high dimensional forecast error covariance estimators. Thus, consistent recovery of the underlying analysis states can be attained by the proposed HD-EnKF procedure.
We establish the mean square error bounds and weak convergence of the analysis states from the HD-EnKF to the underlying analysis under both fixed and high dimensional cases, as well as for both correct and imperfect physical models. The MSE bounds for the multiple-step forecasting and data assimilation are also provided. Simulation studies suggest that the proposed HD-EnKF algorithm could achieve more accurate and robust assimilation results than the widely used inflation-based EnKF method. We apply the proposed HD-EnKF to assimilate sea temperature by combining a physical ocean model with satellite observations and use the ARGO data to validate the quality of the assimilated sea temperature. Both the simulation and the empirical study showed advantages of the HD-EnKF approach.
@@
The paper is structured as follows. Section 2 reviews the Kalman Filter and the EnKF. Section 3 presents several consistent estimators of the high dimensional forecast error covariance and the key Kalman gain matrix, and establishes the theoretical properties of the EnKF. The effects of imperfect or misspecified models on the accuracy of the EnKF are provided in Section 4. Results of empirical data assimilation based on two well-known geophysical models are reported in Section 5 and the application to the sea temperature data assimilation is presented in Section 6, followed by a conclusion in Section 7. Details of the proposed HD-EnKF algorithms, additional theoretical and simulation results, and technical details are given in the Supplementary Material (SM).
@@ 2 Review on Kalman and Ensemble Kalman Filters
where $\mathbf{x}_{t}$ is the $p-$dimensional not fully observed state vector of a dynamic system at time $t$, $\mathbf{M}_{t}$ is the linear model operator (also called evolution or forecast operator), $\mathbf{w}_{t}$ is the model error term with $\mathbf{Q}_{t}$ as its covariance, $\mathbf{y}_{t}$ is the $q-$dimensional observation vector at time $t$, $\mathbf{H}_{t}$ is a linear observation operator which relates the state $\mathbf{x}_{t}$ to the observation $\mathbf{y}_{t}$, and $\boldsymbol{\varepsilon}_{t}$ is the observation error with $\mathbf{R}_{t}$ being its covariance matrix.
@@
In the update (assimilation) step, the goal is to update $\mathbf{x}_{t}^{f}$ according to the new observation $\mathbf{y}_{t}$ and to reduce the difference between the true state $\mathbf{x}_{t}$ and the forecast state $\mathbf{x}_{t}^{f}$. The joint distribution of $\left(\mathbf{x}_{t},\mathbf{y}_{t}\right)$ given $\mathbf{y}_{1:t-1}$ is Gaussian, which implies that
$\mathbf{x}_{t}\mid\mathbf{y}_{1:t}\sim{\mathrm{N}}_{p}\left({\mathbf{x}}_{t}^{a},{\boldsymbol{\Sigma}}_{t}^{a}\right),$ (2.4)
@@
which are, respectively, the analysis (assimilation) state and the associated analysis covariance matrix, and the Kalman gain matrix. Note that the analysis state $\mathbf{x}_{t}^{a}$ is the minimum variance unbiased estimator for the mean of $\mathbf{x}_{t}$ under the linear state-space model (2.1).
However, models in earth science or engineering are often nonlinear, that is the evolution model in (2.1) may not be linear so that $\mathbf{x}_{t}=\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}$, where $\mathcal{M}_{t}$ is a nonlinear model operator that can invalidate the forecast distribution in (2.2). Besides, for large dynamic systems, the dimension $p$ of the state vector $\mathbf{x}_{t}$ can be large so that the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ is expensive to compute and store. To tackle these problems, Evensen (1994) proposed the EnKF, where the forecast error covariance is no longer updated as (2.3) but by the sample covariance based on $n$ forecast ensembles. It hence only needs to store and operate on $n$ vectors of dimension $p$, and is convenient for parallel computing.
@@
where $\mathcal{M}_{t}$ is the possibly nonlinear model operator, $\mathbf{H}_{t}$ is a linear observation operator. It further assumes that the forecast error $\boldsymbol{\eta}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$ follows a Gaussian distribution given $\mathbf{y}_{1:t-1}$, that is $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}_{p}(\mathbf{0},\boldsymbol{\Sigma}_{t}^{f})$, where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)$. Here $\mathcal{M}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are assumed to be known for the time being as in the Kalman Filter.
@@
    Estimate the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance matrix of the forecast ensemble, namely
@@
    and obtain an estimate of the Kalman gain matrix $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$.
@@
The EnKF does not require the expensive recursions (2.3) and (2.5)-(2.6) used in the Kalman Filter. Instead, it estimates the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance. For the case with non-linear model operator as in (2.7), we can still define the underlying analysis state $\mathbf{x}_{t}^{a}$ by (2.5)-(2.6), where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)=\mathbb{E}(\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}\mid\mathbf{y}_{1:t-1})$ is the forecast state and $\boldsymbol{\Sigma}_{t}^{f}$ is the forecast error covariance defined by $\boldsymbol{\Sigma}_{t}^{f}=\operatorname{var}((\mathbf{x}_{t}-\mathbf{x}_{t}^{f})\mid\mathbf{y}_{1:t-1})$. In fact, $\mathbf{x}_{t}^{a}$ defined by (2.5)-(2.6) is the solution to the minimizing problem (2.11)
$J(\mathbf{x})=\frac{1}{2}\big{\{}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})^{\mathrm{T}}(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})+\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)^{\mathrm{T}}\mathbf{R}_{t}^{-1}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)\big{\}}$ (2.11)
with $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ replaced by $\mathbf{x}_{t}^{f}$ and $\boldsymbol{\Sigma}_{t}^{f}$. For the ease of reference, and to avoid differentiating between the linear and the non-linear cases, we also call $\mathbf{x}_{t}^{a}$ as the analysis state of the Kalman Filter even in the non-linear model case. The next two sections will establish the convergence of the EnKF analysis state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman Filter.
3 Convergence of Ensemble Kalman Filter